It's been estimated that if you consume an extra 3500 calories, you will gain 1 pound of a fat. One can of soda contains 40 grams of sugar. Sugar has 4 calories per gram.If a person drinks a six-pack of soda everyday, how many pounds will he or she gain?
Answer by rapaljer(4671)

A can of soda contains 40 grams of sugar which is 40*4= 160 calories. A 6 pack contains 160*6 = 960 calories. If a person drinks a six pack per day, this would be 960 calories *7=6720 calories per week. Divide this by 3500 calories per pound, and you have about 6720/3500=1.92 pounds per week. In a year this would be (mult by 52) about 99.84 pounds ( about 100 pounds per year!!).

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida